OpenStudy (anonymous):

help ?!?! (will give a medal) 1.write an equation of a line parallel to one of the lines in slope-intercept form that contains the point (1,-3). You must show all work and steps to receive credit.(i have chosen line (9,1),(5,7) and line (7,1),(3,7) to be parallel line) 2. write an equation of a line perpendicular to one of the lines in slope-intercept form that contains the point (-2, 7). You must show all work and steps to receive credit. (i have chosen line (1,1),(7,10) and line (9,1),(1,10) are not perpendicular)

OpenStudy (anonymous):

equation for any linear equation is y=mx+b y being your y cordinate. x being ur x. then u got B being the y intercept and m being ur slope. Now slope is just a fancy word for change or how much u gained or lost. When you gain more its a positive, when u lose its a negative. Okay so now lets choose (9,1), (5,7). you can find the slope or change. 7-1=6, 5-9=-4 so the slope is 6/-4, u can simplify this to -(3/2). i believe its safe to assume not looking at your graph that the lines are sloping down. Remeber rise over run. For every 3 units u go to the down u move 2 units to the right. now u figure out the y intercept by looking at the graph given to you and just plug and play baby. i think i need a break on your second question...

OpenStudy (anonymous):

@Studenttrainee for 3. y=-3/2x-3/2

OpenStudy (anonymous):

yes thats it. Good job @luka1998

OpenStudy (anonymous):

@Studenttrainee thx

OpenStudy (anonymous):

@Studenttrainee is it the same way for 4. ??

OpenStudy (anonymous):

@esshotwired

OpenStudy (anonymous):

@amistre64 help plz